OpenStudy (anonymous):

Two redwood trees are the same height. A third tree's height is x feet. The first tree is 50 feet more than three times the height of the third tree. The second tree is 100 feet taller than twice the height of the third tree. What is the height, in feet, of the third tree?

OpenStudy (radar):

x=height of 3rd tree........Problem says so. 3x + 50 = height of 1st tree.......Problem says so. 2x + 100 = height os 2nd tree........Problem says so. 3x + 50 = 2x + 100 Problem says so. (First sentence, two redwood trees are same height. Solve for x using algebra.